4. If the roots of the cubic az3 + bz2 + cz + d = 0 form an arithmetic progression α − β,
α, α + β, prove that (2b
2 − 9ac)b + 27a
2d = 0.
1
Expert's answer
2020-05-20T19:05:52-0400

By the Vieta's formulas we have "-\frac{b}{a}=(\alpha-\beta)+\alpha+(\alpha+\beta)=3\alpha", "\frac{c}{a}=(\alpha-\beta)\alpha+(\alpha-\beta)(\alpha+\beta)+\alpha(\alpha+\beta)="

"=\alpha^2-\alpha\beta+\alpha^2-\beta^2+\alpha^2+\alpha\beta=3\alpha^2-\beta^2", "-\frac{d}{a}=(\alpha-\beta)\alpha(\alpha+\beta)=\alpha^3-\alpha\beta^2"

Then "(2b^2-9ac)b+27a^2d=a^3\left(2\left(\frac{b}{a}\right)^2-9\frac{c}{a}\right)\frac{b}{a}+27a^3\frac{d}{a}"

So we need to prove that "\left(2\left(\frac{b}{a}\right)^2-9\frac{c}{a}\right)\frac{b}{a}+27\frac{d}{a}=0"

We have "2\left(\frac{b}{a}\right)^2-9\frac{c}{a}=2(3\alpha)^2-9(3\alpha^2-\beta^2)=-9\alpha^2+9\beta^2", so "\left(2\left(\frac{b}{a}\right)^2-9\frac{c}{a}\right)\frac{b}{a}+27\frac{d}{a}="

"=(-9\alpha^2+9\beta^2)(-3\alpha)+27(-\alpha^3+\alpha\beta^2)=0"